# Heun-function analysis of the Dirac spinor spectrum in a sine–Gordon soliton background

### v1

@@
We study the Dirac spectrum in a sine–Gordon soliton background, where the induced position-dependent mass reduces the spectral problem to a Heun-type differential equation. Bound and scattering sectors are treated within a unified framework, with spectral data encoded in Wronskians matching local Heun solutions and exhibiting explicit dependence on the soliton parameters and the bare fermion mass. This formulation enables a systematic analysis of spinor bound and scattering states, supported by analytic and numerical verification of wave function matching across the soliton domain. The present work is related to arXiv:2512.07658 and emphasizes a pedagogical treatment of scattering states within the Heun-equation formalism.
@@ 1 Introduction
One of the key strengths of Heun’s equation lies in its ability to capture a wider variety of physical behaviors, including those associated with multi-scale interactions. In such cases, no single asymptotic expansion suffices to describe the system’s behavior across all scales; instead, a global analytic framework is required to connect disparate regions of the solution. Heun’s equation, with its multiple singularities, provides exactly such a global description, offering a unified approach to systems with complex, scale-dependent interactions.
@@
This connection between physical multiplicity (scales, poles, or sources) and analytic multiplicity (singular points) explains why Heun’s equation has become a universal descriptor of complex modern systems — from quantum fields around black holes to molecular and condensed-matter spectra. It captures the interference, coupling, and transition phenomena that occur when multiple characteristic structures coexist in a single system.
In quantum mechanics, Heun’s functions frequently appear in the study of: Anharmonic oscillators and double-well potentials. Two-center Coulomb problems (e.g., diatomic molecules). Quasi-exactly solvable systems, where a finite portion of the spectrum can be obtained analytically. The confluent forms of the Heun equation (e.g., confluent, biconfluent, double confluent) arise naturally when studying potentials such as the Mathieu, Razavy, or double sinh-Gordon potentials.
Heun equations are ubiquitous in modern gravitational and black hole physics. The Teukolsky equation, which governs perturbations of Kerr, Kerr–Newman, or Kerr–de Sitter spacetimes, can often be reduced to confluent or general Heun equations. The radial and angular parts of wave equations for scalar, electromagnetic, or gravitational fields in these geometries are Heun-type. This connection provides a bridge between exact analytical results and numerical relativity, relevant to quasi-normal modes and black hole spectroscopy.
In gauge/gravity duality, AdS/CFT correspondence and string theory, Heun-type equations arise in the analysis of field propagation in AdS black hole backgrounds, fluctuation spectra of D-branes, conformal blocks in 2D CFTs, where Heun functions encode the monodromy structure of conformal correlators. Thus, Heun functions serve as a link between differential geometry, conformal symmetry, and quantum field theory.
In condensed matter and optical physics, Heun equations describe graphene in external fields, photonic crystal modes, quasi-periodic potentials and Bloch waves, topological insulator edge states in nontrivial geometries. Their solutions describe localization, tunneling, and spectral band structures in systems with nonuniform or composite potentials.
@@
From a mathematical physics perspective, Heun’s equation sits at the core of a hierarchy of Fuchsian equations. Its confluent limits define an entire taxonomy of special functions, each relevant to a distinct class of physical systems. Recent progress in symbolic computation (e.g., Maple, Mathematica) and numerical algorithms has made it feasible to use Heun functions in practical calculations, bridging the gap between exact analytic methods and computational physics.
In this paper we study the spectrum of a Dirac fermion in a sine–Gordon soliton background, where the induced position-dependent mass reduces the Dirac equation to a nontrivial eigenvalue problem with spinor components obeying Heun-type equations. This framework resolves the bound-state and scattering spectra in a unified manner, making explicit their dependence on the soliton parameters and bare fermion mass. We show that the Heun formalism is essential for constructing all nonzero-energy states, with the full scattering data encoded in Wronskians matching local Heun solutions across the soliton. The resulting spectral characterization provides a systematic basis for analyzing soliton–fermion stability and topologically protected excitations.
@@ 2 Dirac spinor coupled to a sine-Gordon soliton
Consider the following system of equations [3]
@@
$(\partial_{t}-\partial_{x})\psi_{R}-Me^{-2i\beta\varphi}\,\psi_{L}$ $=$ $0,,$ (2.2)
where $\varphi$ is a sine-Gordon type soliton and $(\psi_{R},\psi_{L})^{T}$ is a two-component Dirac spinor. Let us define the parameterization
@@
where the stationary spinor components $u(x)$ and $v(x)$ define the scattering solutions in the presence of the soliton $\varphi$, and $E_{1}$ is the energy of these states.
@@
We assume the next Ansatz for the scalar field as
$e^{2i\beta\varphi}=-(\frac{1+ie^{-2Kx}}{1-ie^{-2Kx}})^{2},$ (2.8)
with $K\in\hbox{\rm I\kern-1.79993ptR}$. The scalar function $\varphi(x)$ can be written as
$\varphi^{\pm}(x)=-\frac{2}{\beta}\arctan{\Big[e^{2Kx}\Big]},\,\,\,\,K\equiv\pm M.$ (2.9)
This is a kink(anti-kink). These functions exhibit the topological charges
$Q_{k(\bar{k})}$ $=$ $\frac{\beta}{2\pi}(\varphi^{\mp}(+\infty)-\varphi^{\mp}(-\infty)),\,\,\,\,\,k=kink,\,\,\,\bar{k}=antikink$ (2.10)
$=$ $\pm\frac{1}{2}.$ (2.11)
Therefore, the state with topological charge $Q_{kink-top}^{(I)}=+\frac{1}{2}$ is the kink and the state with $Q_{kink-top}^{(I)}=-\frac{1}{2}$ is the anti-kink. Notice that these charges are fractional, i.e. one-half of the integer $\pm 1$. In the Fig. 1 we plot this kink type soliton.
@@
$v(x)=\frac{i}{M}(\frac{1+ie^{-2Kx}}{1-ie^{-2Kx}})^{2}(E_{1}u-iu^{\prime}(x)).$ (2.12)
@@
Figure 1: (color online) Scalar kink (blue) for $M=1.5,\beta=1.$
@@ 2.1 Heun’s equation
In addition, the terms with double poles at $z=a$ become
@@
So, comparing the numerator of (2.40) with the relevant term in (1.1), i.e. $(\alpha\beta z-q)$, one has
@@
$=$ $-\frac{(1+e_{1}e_{2})k^{2}}{4K^{2}}+ie_{1}\frac{(3-e_{1}e_{2})k}{4K}+\frac{iE_{1}}{K}+(1-ie_{1}\frac{k}{K})\sigma,$
@@
$=$ $-\frac{ie_{1}(1+e_{1}e_{2})k}{K}+2\sigma-1.$
The last relationship (2.43) follows from (1.2) and (2.31). Then, one has
@@
In the following we consider the second case $e_{1}e_{2}=-1$ above, i.e. the Eqs. (2.49)-(2.51). Let us consider the next parameter values
$\sigma=0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1.$ (2.52)
Then one has
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=i\frac{E_{1}+e_{1}k}{K}$ (2.53)
$\gamma$ $=$ $\delta^{\star}=1-e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=-e_{1}\frac{ik}{2K}$ (2.54)
@@
where the notation $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is used for the six parameter local Heun function [4, 5]. Note that we have used the parameter set provided in (2.52)-(2.54) for $e_{1}e_{2}=-1,e_{1}=1.$ In fact, the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow ie^{-2Kx}$ in the limit $x\rightarrow+\infty$. So, the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
@@
Remarkably, this equation is similar to the first case equation in (2.23) except that the sign of the term with the factor $(\frac{iE_{1}}{K})$ is reversed. Therefore, in order to write (2.62) in the canonical form (1.1) one performs similar gauge transformation as product of simple powers at the finite singularities as in (2.24)-(2.25). Following similar for the case $e_{1}e_{2}=-1$ above, we can get the next set of parameters
$\sigma$ $=$ $0,\,\,\,e_{3}=1,\,\,\,\,e_{1}=\pm 1.$ (2.63)
$\alpha$ $=$ $-1,\,\,\,\,\beta=0,\,\,\,\,\,q=-i\frac{E_{1}+e_{1}k}{K}$ (2.64)
$\gamma$ $=$ $\delta^{\star}=1+e_{1}\frac{ik}{K},\,\,\,\,\mu=-\nu=e_{1}\frac{ik}{2K}.$ (2.65)
@@
where the parameters were taken from (2.63)-(2.65) with $e_{1}=1$.
@@
Third, if $\gamma\neq\hbox{\sf Z\kern-3.99994ptZ}$, the Heun equation has two linearly independent solutions near $z=0$ [4, 7, 5]. From the relationship (2.24)-(2.25), since $\frac{k}{K}$ is real for the scattering states, one can get two linearly independent solutions $u_{1}^{(1)}(z)$ and $u_{1}^{(2)}(z)$ as
@@
where the parameters (2.52)-(2.54) can be used.
A similar construction can be performed for the second case, providing two linearly independent solutions $u_{2}^{(1)}(z)$ and $u_{2}^{(2)}(z)$ with parameters (2.63)-(2.65). In fact, for the second type of solutions the argument of the local Heun function tend to zero as $x\rightarrow$-$\infty$, meaning that the local Heun function remains analytic in this limit. Note that the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow-ie^{2Kx}$ from (2.67). Then, a first solution of (2.62) from (2.66)-(2.67) can be written as
$u_{2}^{(1)}(x)$ $=$ $z^{\mu}(z-1)^{\nu}\,Hl[\frac{1}{2},-i\frac{E+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}],$ (2.71)
$=$ $e^{-\frac{\pi k}{4K}}e^{ikx}Hl[\frac{1}{2},-i\frac{E+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}]$ (2.72)
where the parameters correspond to the set (2.63)-(2.65) with $e_{1}=1$.

### v2

@@
We study the Dirac spectrum in a sine-Gordon soliton background, where the induced position-dependent mass reduces the spectral problem to a Heun-type differential equation. Bound and scattering sectors are treated within a unified framework, with spectral data encoded in Wronskians matching local Heun solutions and exhibiting explicit dependence on the soliton parameters and the bare fermion mass. This formulation enables a systematic analysis of spinor bound and scattering states, supported by analytic and numerical verification of wave function matching across the soliton domain. The present work is related to arXiv:2512.07658 and emphasizes a pedagogical treatment of scattering states within the Heun-equation formalism.
@@ 1 Introduction
Topological soliton–fermion systems lead naturally to spectral problems governed by second-order differential equations with multiple singularities. In the presence of a sine–Gordon soliton background, the Dirac equation reduces to a Fuchsian system that can be systematically mapped into the canonical Heun equation. This structure reflects the underlying multi-scale and topological features of the interaction and provides a unified analytic setting for bound and scattering states.
@@
In this paper we study the spectrum of a Dirac fermion in a sine–Gordon soliton background, where the induced position-dependent mass reduces the Dirac equation to a nontrivial eigenvalue problem with spinor components obeying Heun-type equations. This framework resolves the bound-state and scattering spectra in a unified manner, making explicit their dependence on the soliton parameters and bare fermion mass. We show that the Heun formalism is essential for constructing all nonzero-energy states (valence fermions), with the full scattering data encoded in Wronskians matching local Heun solutions across the soliton. The resulting spectral characterization provides a systematic basis for analyzing soliton–fermion stability and topologically protected excitations.
The present work builds upon Ref. [3] by providing a pedagogical treatment of the transformation of Fuchsian equations into the canonical Heun form. Two distinct mapping procedures are introduced, and a systematic analysis of additional parameter sets is carried out to clarify the structure and equivalence of the resulting formulations. We explain qualitatively (through additional plots not presented in [3]) the incident, reflected and transmitted scattering waves, as well as the phase shifts for a set of fermion mass parameters. The physical interpretation of Dirac current conservation, and its equivalence to the topological current of the scalar field, is examined within the framework of scattering unitarity. New results regarding the Fuchsian $\rightarrow$ Heun transformation procedure are presented and discussed in the Appendices A and B.1.
@@ 2 Dirac spinor coupled to a sine-Gordon soliton
Consider the following system of equations [^1] Our notation: $\partial_{0}=\frac{\partial}{\partial t},\partial_{1}=\frac{\partial}{\partial x}$, $\not{\partial}=\gamma^{\mu}\partial_{\mu}$. We use $\gamma_{0}=\left(\begin{array}[]{cc}0&i\\ -i&0\end{array}\right)$, $\gamma_{1}=\left(\begin{array}[]{cc}0&-i\\ -i&0\end{array}\right)$, $\gamma_{5}=\gamma_{0}\gamma_{1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, and $\psi=\left(\begin{array}[]{c}\psi_{R}\\ \psi_{L}\end{array}\right),\,\,\bar{\psi}=\psi^{\dagger}\gamma_{0},\,\psi_{R}\equiv(\frac{1+\gamma_{5}}{2})\psi,\,\psi_{L}\equiv(\frac{1-\gamma_{5}}{2})\psi$. [3, 2]
$i\not{\partial}\psi=Me^{2i\beta\varphi\gamma_{5}}\psi,$ (2.1)
and
$j^{\mu}$ $=$ $-\frac{1}{\beta}\,\epsilon^{\mu\nu}\partial_{\nu}\varphi,\,\,\,\,\,\,\,j^{\mu}\equiv\bar{\psi}\gamma^{\mu}\psi,\,\,\,\,\mu=0,1,\,\,\nu=0,1,\,\,\,\,\epsilon^{01}=-\epsilon^{10}=1,\epsilon^{00}=\epsilon^{11}=0.$ (2.2)
In components the Eq. (2.1) becomes
@@
$(\partial_{t}-\partial_{x})\psi_{R}-Me^{-2i\beta\varphi}\,\psi_{L}$ $=$ $0.$ (2.4)
Whereas, the equation (2.2) in components becomes
$j^{0}$ $=$ $\psi^{\star}_{R}\psi_{R}+\psi^{\star}_{L}\psi_{L}=-\frac{1}{\beta}\,\partial_{x}\varphi,$ (2.5)
$j^{1}$ $=$ $-\psi^{\star}_{R}\psi_{R}+\psi^{\star}_{L}\psi_{L}=-\frac{1}{\beta}\,\partial_{t}\varphi.$ (2.6)
Equation (2.2) establishes an equivalence between the Noether current $j^{\mu}$ associated with the Dirac spinor and the topological current $\epsilon^{\mu\nu}\partial_{\nu}\varphi$ of the real scalar field. This equivalence is a key structural feature of certain integrable models, notably affine Toda models coupled to fermions [2].
2.1 Spinor zero-mode coupled to a soliton
Let us find a static solution of the above fermion-soliton system. We assume the next Ansatz for the scalar field as
$e^{2i\beta\varphi}=-\left(\frac{\tau}{\tau^{*}}\right)^{2},\,\,\,\,\tau=1+ie^{-2Kx},\,\,\,\,\,K\in\hbox{\rm I\kern-1.79993ptR}.$ (2.7)
The real scalar function $\varphi(x)$ is defined by taking the $log^{\prime\prime}$ function of (2.7) and considering its principal branch as $i\beta\varphi=\mbox{log}{|-(\frac{\tau}{\tau^{*}})^{2}|}+i\arg{(-(\frac{\tau}{\tau^{*}})^{2})}$. The function $\varphi$ can be written as
$\varphi=\frac{2}{\beta}\arctan{[\frac{1}{i}\,\frac{\tau-\tau^{\star}}{\tau+\tau^{\star}}]}\pm\frac{\pi}{2\beta}.$ (2.8)
Then, one has
$\varphi(x)=\frac{2}{\beta}\arctan{\Big[e^{-2Kx}\Big]}\pm\frac{\pi}{2\beta}.$ (2.9)
This is a kink(anti-kink). Note that the asymptotic values become: 1) $K<0$ one has $\varphi(+\infty)=\frac{\pi}{\beta}\pm\frac{\pi}{2\beta}$ and $\varphi(-\infty)=\pm\frac{\pi}{2\beta}$. 2) $K>0$ one has $\varphi(+\infty)=\pm\frac{\pi}{2\beta}$ and $\varphi(-\infty)=\frac{\pi}{\beta}\pm\frac{\pi}{2\beta}$.
These functions exhibit the topological charges
$Q_{k(\bar{k})}$ $=$ $\frac{\beta}{2\pi}(\varphi(+\infty)-\varphi(-\infty)),\,\,\,\,\,k=kink\,(K<0),\,\,\,\bar{k}=antikink\,(K>0)$ (2.10)
$=$ $\pm\frac{1}{2}.$ (2.11)
Therefore, the state with topological charge $Q_{kink-top}^{(I)}=+\frac{1}{2}$ is the kink and the state with $Q_{kink-top}^{(I)}=-\frac{1}{2}$ is the anti-kink. Notice that these charges are fractional, i.e. one-half of the integer $\pm 1$. In the Fig. 1 we plot this kink type soliton for $K=-M,\,\,M=\frac{3}{2},\,\,\beta=1$ and with the minus sign in the shifting constant in (2.9), i.e. ($\frac{\pi}{2\beta}$). The zero-mode spinor components become
$\left(\begin{array}[]{c}\psi_{R}\\ \psi_{L}\end{array}\right)$ $=$ $c_{1}\left(\begin{array}[]{c}\frac{\zeta\,e^{-Kx}}{1-ie^{-2Kx}}\\ -\frac{e^{-Kx}}{1+ie^{-2Kx}}\end{array}\right),\,\,\,\,\ \zeta=\mbox{sign}(\frac{K}{M}),\,\,\,K=\pm M.$ (2.16)
The static kink (2.9) and spinor (2.16) are solutions of the firt order system of nonlinear equations (2.3)-(2.4) and (2.5)-(2.6).
Let us emphasize that the system of soliton-fermion system (2.1)-(2.2) possesses a kink type solution (2.9) with fractional topological charge (2.10) coupled to the spinor bound state (2.16). The parameter relationship $K=\pm M$ must hold in order to have this soliton-fermion solution. One can see that the soliton’s width becomes $\sim\frac{1}{2M}$, whereas the spinor bound state charge density $j^{0}$ is confined inside the soliton (see Fig.1). This solution can be considered to be a zero-mode bound state, as we will explain below.
2.2 Spinor bound states and scattering modes coupled to a soliton
Here we consider the spinor bound states (non-zero modes ou valence fermions) and scattering states coupled to the soliton (2.9) written in the form (2.7). So, let us define the parameterization
@@ 2.2 Spinor bound states and scattering modes coupled to a soliton
where the stationary spinor components $u(x)$ and $v(x)$ define the bound states or the scattering solutions in the presence of the soliton $\varphi$, and $E_{1}$ is the energy of these states.
@@
$v(x)=\frac{i}{M}(\frac{1+ie^{-2Kx}}{1-ie^{-2Kx}})^{2}(E_{1}u-iu^{\prime}(x)),$ (2.22)
where the exponential $e^{2i\beta\varphi}$ has been substituted using the relationship (2.7).
@@
Figure 1: (color online) Scalar kink (blue) and spinor charge density (red) for $M=1.5,\beta=1.$
Note that the system (2.20)-(2.21) can be written as an eigenvector equation
$H\left(\begin{array}[]{c}u(x)\\ v(x)\end{array}\right)=E_{1}\left(\begin{array}[]{c}u(x)\\ v(x)\end{array}\right),\,\,\,\,\,H\equiv\left(\begin{array}[]{cc}i\frac{d}{dx}&iMe^{-2i\beta\varphi(x)}\\ -iMe^{2i\beta\varphi(x)}&-i\frac{d}{dx}\end{array}\right).$ (2.30)
Note that the Hamiltonian matrix $H$ for real scalar field $\varphi$ is a Hermitian operator, i.e. $H^{\dagger}=H$, for $\beta$ and $M$ real parameters. The real scalar field condition implies the spinor current density $j^{\mu}$ to be real in the equation (2.2). In addition, one can show that the current is conserved. In fact, by taking the $\mu$-derivative of (2.2) one can show
$\partial_{\mu}j^{\mu}=0,$ (2.31)
where the completely antisymmetric property of $\epsilon^{\mu\nu}$ has been used. Therefore from (2.31) one can show that the charge $Q=\int_{-\infty}^{+\infty}dx\,j^{0}$ is a conserved quantity. This is related to the topological charge, as one can see from (2.2) by integrating in the whole real $x$-coordinate its 00-component
$Q$ $=$ $\int_{-\infty}^{+\infty}dx\,j^{0}$ (2.32)
$=$ $-\frac{1}{\beta}(\varphi(x=+\infty)-\varphi(x=-\infty))$
$=$ $-\frac{2\pi}{\beta^{2}}Q_{k(\bar{k})},$
where in the last equation it has been used the relationship (2.10).
Some comments are in order here.
First, the solution (2.16) becomes the zero-mode bound state sector of the system of equations (2.20)-(2.21), since for $E_{1}=0$ these equations become the static versions of the system (2.3)-(2.4) with the identifications $\psi_{R}(x,0)=u(x)$ and $\psi_{L}(x,0)=v(x)$ consistent with the Ansatz (2.19).
Second, for the Dirac spinor interacting with the kink unitarity follows from the conservation of the Dirac current (2.31) provided that the Hamiltonian is hermitian.
Third, at spatial infinity the solutions reduce to free Dirac plane waves, as can be seen in the system (2.20)-(2.21) for the scalar soliton field approaching non-vanishing constant values in the asymptotic regions $x\rightarrow\pm\infty$. In the Appendix B.1 we discuss the asymptotic scattering and bound states.
Fourth, unitarity implies conservation of the spatial component of the Dirac current, $j^{1}_{in}=j^{1}_{ref}+j^{1}_{tr}$, which amounts to the relation $R+T=1$, with reflection and transmission coefficients defined using the flux, i.e. $R=\frac{j^{1}_{ref}}{j^{1}_{in}}$ and $T=\frac{j^{1}_{tr}}{j^{1}_{in}}$.
Fifth, in our case unitarity is tied to the topological structure of the theory due to the equation (2.2). The fermionic probability flux is conserved as long as the topological charge associated with $\varphi$ is conserved. Reflection and transmission can be interpreted as redistribution of topological flux rather than loss or gain of probability.
@@ 2.3 Heun’s equation
In addition, the terms with double poles at $z=a$ in (2.53)-(2.53) become
@@
So, comparing the fraction (2.59) with the relevant factor in (1.1), i.e. $\frac{\alpha\beta z-q}{z(z-1)(z-a)}$, one has
@@
$=$ $-\frac{(1+e_{1}e_{2})k^{2}}{4K^{2}}+ie_{1}\frac{(3-e_{1}e_{2})k}{4K}+\frac{iE_{1}}{K}+(1-ie_{1}\frac{k}{K})\sigma.$
In addition, from (1.2) and (2.50) one has
@@
$=$ $-\frac{ie_{1}(1+e_{1}e_{2})k}{K}+2\sigma-1.$ (2.63)
Then, from the above relationships one has
@@
In the Appendix B.2, Eqs. (B.6)-(B.8) we consider the second case $e_{1}e_{2}=-1$ above, i.e. the Eqs. (2.69)-(2.71).
@@
where the notation $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is used for the six parameter local Heun function [4, 5]. Note that we have used the parameter set provided in (B.6)-(B.8) for $e_{1}e_{2}=-1,e_{1}=1.$ In fact, the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow ie^{-2Kx}$ in the limit $x\rightarrow+\infty$. So, the argument of the local Heun function tend to zero (one of the regular singular points at $z=0$) as $x\rightarrow+\infty$. The local Heun function $Hl[\frac{1}{2},q;\alpha,\beta,\gamma,\delta;z]$ is analytic at the regular singular point $z=0$ and is normalized such that $Hl(0)=1$. Around this point, it admits a convergent Taylor expansion. In the complex $z$-plane the radius of convergence of this local series is given by $min\{|a|,1\}$; for $a=1/2$, this yields a convergence radius of $1/2$.
@@
Remarkably, this equation is similar to the first case equation in (2.42) except that the sign of the term with the factor $(\frac{iE_{1}}{K})$ is reversed. Therefore, in order to write (2.79) in the canonical form (1.1) one performs similar gauge transformation as product of simple powers at the finite singularities as in (2.43)-(2.44). Following similar steps for the case $e_{1}e_{2}=-1$ above, one can get the relevant parameters. In the Appendix B.3, Eqs. (B.11)-(B.13) we provide the corresponding parameters.
@@
where the parameters were taken from (B.11)-(B.13) with $e_{1}=1$.
@@
[…]
@@
where the parameters (B.6)-(B.8) can be used.
A similar construction can be performed for the second case, providing two linearly independent solutions $u_{2}^{(1)}(z)$ and $u_{2}^{(2)}(z)$ with parameters (B.11)-(B.13). In fact, for the second type of solutions the argument of the local Heun function tend to zero as $x\rightarrow$-$\infty$, meaning that the local Heun function remains analytic in this limit. Note that the behavior of $z$ in the original spatial coordinate becomes $z\rightarrow-ie^{2Kx}$ from (2.81). Then, a first solution of (2.79) from (2.80)-(2.81) can be written as
$u_{2}^{(1)}(x)$ $=$ $z^{\mu}(z-1)^{\nu}\,Hl[\frac{1}{2},-i\frac{E_{1}+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}],$ (2.85)
$=$ $e^{-\frac{\pi k}{4K}}e^{ikx}Hl[\frac{1}{2},-i\frac{E_{1}+k}{K};-1,0,1+i\frac{k}{K},1-i\frac{k}{K};\frac{1}{1+ie^{-2Kx}}]$ (2.86)
where the parameters correspond to the set (B.11)-(B.13) with $e_{1}=1$.